# Mix-of-Granularity: Optimize the Chunking Granularity for Retrieval-Augmented Generation

### v1

@@ 3.2.2 Soft-labels
In MoG, we use a Multi-Layer Perceptron (MLP) as the router, which is trained in a supervised learning method. The input for the router is the embedding of the input query $q$ generated with RoBERTa [32], which is then mapped to $w$ by the router. The intrinsic training signal is the “labels” ($l$) in the Medical Question-Answering (MQA) datasets. In some MQA datasets, the ground truth snippets to be retrieved are provided, so we use them as “labels” directly; otherwise, the concatenation of the strings of the “Question” and “Answer” is used as the “label”. A natural training objective is to maximize the semantic similarity between $chunk_{r}$ and $l$ by adjusting $w$. Unfortunately, this label can not be used to guide the training directly because there is a non-differentiable top-$k$ selection in the way. The soft labels are proposed to bypass the top-$k$ selection during the training.
For each query $q$, the most relevant snippet ($S_{best}$) is retrieved from the reference documents of each granularity level. The semantic similarity between each snippet in $S_{best}$ and the label $l$ is then calculated (with static models including TF-IDF [38], RoBERTa [32], or hitrate score) and stored in $sim_{best}$. We create a soft label of 0.8 (resp. 0.2) for the most (resp. the second) similar snippet in $S_{best}$, and pad 0 for the other snippets.
With the soft labels ($sl$) built, we can train the router to predict a high value (0.8) for the optimal granularity level, while conserving certain flexibility to choose the second-best granularity level. The router is trained by minimizing a Binary Cross Entropy loss function ($l_{bce}$):
@@ 3.3 MoGG: MoG with Graph-context
In MoG, by adjusting the granularity level, the adjacent snippets with relevant knowledge can be retrieved altogether. This method is particularly effective when information centered around the same topic is stored in sentences close to each other. However, in most cases, answering a complex question requires reasoning over information stored in different paragraphs or even different documents. In practice, the solution is to perform more retrievals at a finer granularity level, retrieving only highly relevant small pieces to form a comprehensive reasoning chain. This approach is inconvenient because determining the optimal number of retrievals often involves manually adjusting $k$ for the final selection, which is challenging because the whole tuning process is time-consuming and $k$ can not go infinitely large. In this scenario, MoG shows no clear advantage compared to classic RAG implementations, because when changing the $k$ is hard, adjusting the granularity level will only sacrifice either the precision or recall of the retrieval.
To overcome this challenge, an intuitive approach is to reorganize the reference document and group the relevant information so that each snippet in Figure 2 contains both the originally dispersed relevant text and the adjacent content. In this way, even for a limited $k$, the selected snippets can encompass information from different paragraphs or documents. Motivated by this, we propose a more applicable framework, MoGG, by enhancing MoG with a preprocessing step that organizes the reference database $\mathcal{K}$ into a graph. As illustrated in Figure 3, each document is initially split into small pieces consisting of one or two sentences, and each piece is treated as a separate node in the graph. To determine the edges in the graph, an index is first created with all these nodes. Then, each node is used as a “query” to search for the $k_{graph}$ (set as 3 in Figure 3) most relevant nodes using a retriever. An edge is then added between the query node and each of the relevant nodes while ensuring that the similarity between them meets a predefined threshold $T_{graph}$. For simplicity, here we use the same retriever as the RAG system. The graph obtained is a binary undirected graph, meaning that the connections between any nodes are bi-directional.
@@
This transformation of the original reference document into a graph extends the concept of “context” in linear text to “neighbors” in a graph. In a graph setting, granularity levels are adapted to “hopping ranges”: in a non-graph setting, a larger granularity level corresponds to more adjacent sentences being grouped in a chunk; in a graph setting, a larger granularity level corresponds to sentences within a larger hopping range of a centered sentence being grouped as one chunk. To avoid context redundancy, duplicate nodes in the neighbors are considered only once. Similar to MoG, the documents in the external knowledge database $\mathcal{K}$ are chunked with $n_{gra}$ different hopping ranges and then encoded into embeddings. The rest of the MoG, i.e., parts I and III in Figure 2, remains unchanged. By transforming the documents into a graph structure and defining granularity based on hopping ranges, MoGG effectively captures dispersed relevant information, allowing for more comprehensive and efficient retrieval.
@@ 4.1 Corpus and Medical QA datasets
A reference knowledge database used in a RAG system is often termed a “corpus”. To form the corpora, data from various sources were collected, including the widely-used PubMed  [^1] https://pubmed.ncbi.nlm.nih.gov/ corpus for all biomedical abstracts, the StatPearls  [^2] https://www.statpearls.com/ corpus for clinical decision support, the Textbooks[21] corpus covering medical textbook knowledge, and the Wikipedia  [^3] https://www.wikipedia.org corpus for general knowledge. The four corpora are combined to form a larger corpus, named “MedCorp”. The statistics of the used corpora are shown in Table 1 (left).
Following the setup in the MIRAGE benchmark [51], we evaluate the performance of the RAG system using five Medical Question-Answering datasets: MMLU-Med [15], MedQA-US [53], MedMCQA [35], PubMedQA* [22], and BioASQ-Y/N [47]. Specifically, we extracted only the multiple-choice questions relevant to the biomedical field and removed the ground truth supporting contexts during testing. The details of each QA dataset are included in Appendix A. The statistics of the five datasets used in our experiments are listed in Table 1 (right).
Table 1: Important statistics of corpora and QA datasets
Corpus #Doc #Snip. $\bar{L}$. Domain
PubMed 23.9M 23.9M 196 Bio.-Med.
StatPearls 9.3k 301.2k 119 Clinics
Textbooks 18 125.8k 182 Medicine
Wikipedia 6.5M 29.9M 162 General
MedCorp 30.4M 54.2M 221 Mixed
Datasets Size #Opt. $\bar{L}$ Source
MMLU-Med 1,089 4 63 Exam.
MedQA-US 1,273 4 177 Exam.
MedMCQA 4,183 4 26 Exam.
PubMedQA* 500 3 24 Literature
BioASQ-Y/N 618 2 17 Literature
@@ 4.2 Experiment Setup
All backbone LLMs, whether accessed via API or local deployment, are run under off-the-shelf settings. The exact versions of the backbone LLMs are listed in Appendix B. Experiments are conducted on Nvidia GeForce 3090 and 4090 GPUs. The code is written using the PyTorch framework, utilizing an Adam optimizer with a learning rate of 0.001. Each training job is run until the convergence of the loss value. During the experiment, there are two top-k selections. Unless specified, when retrieving snippets from each corpus, we select the top-3 snippets; when all the snippets are retrieved from each corpus, we select the top-2 snippets with the highest relevance scores to pass to the backbone LLM. When snippets are too long, they are truncated automatically to fit the LLM’s context window size. The router requires approximately 12GB of GPU memory for training and 6GB for inference. Utilizing a caching mechanism, we efficiently completed 35 training sessions and over a hundred inferences, each training session taking around 4 hours for 1000 epochs.
4.3 Performance of MoG on Medical QA Task
As mentioned above, we test the effectiveness of MoG on Medical Question-Answering datasets. Each question within the dataset is formulated as a multiple-choice question. The LLM is tasked with choosing the best answer option(s) from the given options. Under this setting, we evaluated the performance of the entire RAG system by calculating the Exact Matching accuracy of the answers. To prevent knowledge leakage, following previous work [51], only the question is used (options not given) to retrieve reference documents from the external knowledge database.
The router of MoG guides the retrieval system to choose the optimal granularity. As shown in Figure 4, when tested on different datasets, the router trained with Textbook corpus shows a preference for different granularity levels. For instance, on the PubMedQA dataset, the finest granularity snippets are selected most frequently. This is because the questions in PubMedQA are typically precise and can be answered with short reference snippets.
@@ 4.3 Performance of MoG on Medical QA Task
MoG is integrated into one same RAG system, with which we conduct the Medical Question-Answering task. Backbone LLMs are altered to cover some of the popular ones, such as ChatGPT [4], InternLM2 [5], Llama3 [1], GLM3 [12], and Qwen1.5 [2]. The router is trained with the soft labels built using RoBERTa [32], this choice is justified by the experiment in Appendix C. Three candidate snippets are retrieved from each granularity level of the external knowledge database. To investigate the effect of varying the number of candidate snippets on RAG system performance, we conducted experiments with different snippet counts. Detailed methodology and results are provided in Appendix D. The retriever is fixed as BM25 [39], with a further discussion on the performance of different retrievers included in Appendix E. Here we present the experimental results obtained with the router trained with MedCorp corpus in Table 2, the results obtained with routers trained on four single corpora are presented in Appendix F. The results demonstrate that MoG consistently enhanced the performance of the RAG system across different backbone models. Apart from the standard baseline using Chain-of-Thought (CoT) [50], we add a RAG system with a single granularity corpus as the baseline. The granularity level selected for this baseline is set to be the same as the one in MedRAG [51], therefore, it is also noted as “MedRAG” in Table 2.
From the table, we find that MoG constantly performs better than MedRAG, while not necessarily better than CoT. The reason why CoT might be better is that the RAG system we used has no noise filters, thus the noise is injected together with the external knowledge via prompts. A detailed analysis of the number of samples improved or degraded by the application of MoG is included in the Appendix G, in which we manually verified that the majority of degradation is caused by noise.
We also find that MoG improves the accuracy score more when applied on smaller, weaker LLMs (like ChatGLM and Qwen), this is probably because smaller LLMs have less knowledge stored in their internal parameters and, thus could benefit more from the retrieved snippets.
Table 2: Accuracy of Medical Question-Answering task with MoG (trained with MedCorp), best results marked in bold.
@@
4.4 Performance of MoGG on Medical QA Task
Similarly to the experiment of MoG, we test the performance of MoGG on Medical QA datasets. Due to the large volume of Wikipedia and PubMed corpus, building the index for them in the setting of MoGG is too time-consuming. Therefore, we tested only the performance of MoGG with routers trained on Textbooks and StatPearls corpora. In Table 3 we present the results obtained when trained with Textbook corpus. The results obtained with StatPearls show similar patterns (listed in Appendix H). From the table, we can tell that MoGG can further improve the averaged accuracy scores. By comparing Table 2 and Table 3, we can find that, even when trained with significantly fewer samples (Textbooks corpus is a tiny subset of MedCorp corpus, accounting for only about 0.2% of all the snippets in MedCorp), MoGG brings more significant improvement in terms of the averaged accuracy score with respect to MedRAG than MoG. This finding highlights that MoGG is more efficient than MoG thanks to its flexible way of organizing the reference snippets (in the form of a graph).
Table 3: Accuracy of Medical Question-Answering task with MoGG (trained with Textbooks), best results marked in bold.
@@ 5 Limitations and Broader Impacts
Our work serves as an early trial in dynamic chunking strategy, with the following major directions for improvement. First, MoG(G)’s candidate granularity levels are manually assigned. It could be more efficient if an algorithm automatically set these granularity levels to avoid excessive grid-searching for parameter optimization. Secondly, the current router uses only the semantic information of the input query to predict the best granularity level. In practice, other factors might influence this choice, such as the user’s knowledge of relevant fields, the type of query, and the expected response length. Incorporating more information into the router could potentially improve prediction results. Lastly, MoG(G) could be combined with other techniques to further enhance retrieval quality, such as Recursive Character Splitting  [^4] https://python.langchain.com/v0.1/docs/modules/data_connection/document_transformers/recursive_text_splitter, Parent Document Retrieval [46], or Sliding Window Chunking [40].
MoG(G) provides a valuable tool for integrating multiple sources of external knowledge databases into a RAG system, thereby enhancing RAG systems with existing knowledge sources. However, the router also introduces a new security risk: a compromised router could redirect knowledge retrieval to malicious sources, injecting incorrect or even harmful information into the backbone LLM. Therefore, it is crucial to protect and monitor the router to mitigate this risk.
@@ 6 Conclusion
In this work, we present Mix-of-Granularity (MoG), a mechanism to dynamically choose the best granularity when retrieving information from an external knowledge database. When applied to a Retrieval-Augmented Generation system, MoG helps retrieve more relevant information while reducing noise. MoG is further extended as Mix-of-Granularity-Graph (MoGG), where reference documents are reorganized in graph form. This extension allows distantly situated information to be retrieved simultaneously, overcoming the limitations of a fixed top-$k$ selection strategy. Finally, we introduce a soft label guided loss function to address the difficulty of backward propagation with top-$k$ selection, which could benefit future research.
Appendix A Details of QA Datasets
A.1 MMLU-Med
The Massive Multitask Language Understanding (MMLU) benchmark [15] evaluates the multitask learning capability of language models. While the full MMLU dataset encompasses 57 different tasks, we specifically extracted the medical questions for our tests, totaling 1089 questions.
A.2 MedQA-US
MedQA [53] is a multiple-choice QA dataset derived from professional medical board exams. It is available in Simplified Chinese, Traditional Chinese, and English. For our experiments, we used 1273 questions from the English version.
A.3 MedMCQA
MedMCQA [35] comprises a large number of questions from the Indian medical entrance exam, covering 2400 healthcare topics and 21 medical subjects. For the MIRAGE benchmark, we utilized the “dev” set of the original MedMCQA dataset.
A.4 PubMedQA*

### v2

@@ 3.2.2 Soft-labels
In MoG, we train a Multi-Layer Perceptron (MLP) as the router in a supervised learning method. The input for the router is the embedding of the input query $q$ generated with RoBERTa Liu et al. (2019), which is then mapped to $w$ by the router. The intrinsic training signal is the “labels” ($l$) in the Medical Question-Answering (MQA) datasets. In some MQA datasets, the ground truth snippets to be retrieved are provided, so we use them as “labels” directly; otherwise, the concatenation of the strings of the “Question” and “Answer” is used as the “label”. A natural training objective is to maximize the semantic similarity between $chunk_{r}$ and $l$ by adjusting $w$. Such a label helps us to train the router to choose the best granularity based on the input query, which is a prior, with solid posterior information (the similarity between the ans and the text of different granularities). Unfortunately, this label can not be used to guide the training directly because there is a non-differentiable top-$k$ selection in the way. The soft labels are proposed to bypass the top-$k$ selection during the training: For each query $q$, the most relevant snippet ($S_{best}$) is retrieved from the reference documents of each granularity level with BM25 Robertson and Zaragoza (2009). The semantic similarity between each snippet in $S_{best}$ and the label $l$ is then calculated (with static models including TF-IDF Ramos (2003), RoBERTa Liu et al. (2019), or hitrate score) and stored in $sim_{best}$. We create a soft label of 0.8 (resp. 0.2) for the most (resp. the second) similar snippet in $S_{best}$, and pad 0 for the other snippets.
For example, the soft labels corresponding to $sim_{best,1}$ [0, 0.32, 0.11, 0.88, 0.45] and $sim_{best,2}$ [0.95, 0.07, 0.22, 0.11, 0.19] are $sl_{1}$ [0, 0, 0, 0.8, 0.2] and $sl_{2}$ [0.8, 0, 0.2, 0, 0], respectively.
The values of the soft labels are designed to guide the router in distinguishing the relative importance among the granularity candidates. Empirically, setting these values to either [0.8, 0.2, 0] or [0.7, 0.3, 0] yields similar results. With the soft labels ($sl$) built, we can train the router to predict a high value (0.8) for the optimal granularity level, while conserving certain flexibility to choose the second-best granularity level. The router is trained by minimizing a Binary Cross Entropy loss function ($l_{bce}$):
@@ 3.3 MoGG: MoG with Graph-context
With MoG, the adjacent snippets with relevant knowledge can be retrieved altogether by adjusting the granularity level. This method is particularly effective when information centered around the same topic is stored in adjoining sentences. However, in most cases, answering a complex question requires reasoning over information stored in different paragraphs or even different documents. A common solution is to perform more retrievals at a finer granularity level, retrieving only highly relevant small pieces to form a comprehensive reasoning chain. This approach is inconvenient because determining the optimal number of retrievals often involves manually adjusting $k$, which is challenging because the whole tuning process is time-consuming and $k$ can not go infinitely large.
To overcome this challenge, an intuitive approach is to reorganize the reference document and group the relevant information together. Motivated by this, we propose a more applicable framework, MoGG, by enhancing MoG with a preprocessing step that organizes the documents in $\mathcal{K}$ as a graph. As illustrated in Figure 3, each document is initially split into small pieces consisting of one or two sentences, and each piece is treated as a separate node in the graph. To determine the edges in the graph, an index is first created with all these nodes. Then, each node is used as a “query” to search for the $k_{graph}$ (set as 3 in Figure 3) most relevant nodes using BM25. An edge is then added between two nodes if the similarity between them meets a predefined threshold $T_{graph}$.
@@
This pre-processing extends the concept of “context” in linear text to “neighbors” in a graph. Also, granularity levels are adapted to “hopping ranges”: in a non-graph setting, a larger granularity level corresponds to more adjacent sentences being grouped in a chunk; in a graph setting, a larger granularity level corresponds to nodes within a larger hopping range of a centered node being grouped as one chunk. To avoid context redundancy, duplicate nodes in the neighbors are considered only once. Similar to MoG, the documents in the external knowledge database $\mathcal{K}$ are chunked with $n_{gra}$ different hopping ranges and then encoded into embeddings, while the rest of the MoG remains unchanged. By transforming the documents into a graph and defining granularity based on hopping ranges, MoGG effectively captures dispersed relevant information, allowing for more comprehensive and efficient retrieval.
@@ 4.1 Corpus and Medical QA datasets
A reference knowledge database used in a RAG system is often termed a “corpus”. To form the corpora, data from various sources were collected, including the widely-used PubMed Galileo Mark Namata and Huang (2012) corpus for all biomedical abstracts, the StatPearls Publishing (2024) corpus for clinical decision support, the TextbooksJin et al. (2020) corpus covering medical textbook knowledge, and the Wikipedia Wikimedia Foundation (2024) corpus for general knowledge. The four corpora are combined to form a larger corpus, named “MedCorp”. For each corpus used in our experiments, there are 5 granularity levels tested. The chunking size of the second granularity level is set to be the same as the one in MedRAG to facilitate the comparison of the results. The chunking size of the {1st, 3rd, 4th, 5th} granularity level is set to be {1/2, 2, 4, 8} times of the one in MedRAG.
The performance of the RAG system are evaluated using five MQA datasets following MIRAGE benchmark Xiong et al. (2024): MMLU-Med Hendrycks et al. (2021), MedQA-US Zhang et al. (2018), MedMCQA Pal et al. (2022), PubMedQA* Jin et al. (2019), and BioASQ-Y/N Tsatsaronis et al. (2015). Specifically, only the biomedical questions are kept and the ground truth supporting contexts are removed during testing.
The detailed descriptions of each copora and each QA dataset, as well as their important statistics, are included in Appendix B. We have begun to test our method using medical question-answering datasets, as they represent a knowledge-intensive domain. We posit that significant improvements demonstrated by the tests on this knowledge-intensive field suggest MoG’s potential effectiveness in other domains with lower knowledge dependencies and higher error tolerances.
@@ 4.2 Experiment Setup
All backbone LLMs, whether accessed via API or local deployment, are run under off-the-shelf settings. The exact versions of the backbone LLMs are listed in Appendix C. Experiments are conducted on Nvidia GeForce 3090 and 4090 GPUs. The code is written using the PyTorch framework, utilizing an Adam optimizer with a learning rate of 0.001. Each training job is run until the convergence of the loss value. During the experiment, there are two top-k selections. Unless specified, when retrieving snippets from each corpus, we select the top-3 snippets; when all the snippets are retrieved from each corpus, we select the top-2 snippets with the highest relevance scores to pass to the backbone LLM. When snippets are too long, they are truncated automatically to fit the LLM’s context window size. The router requires approximately 12GB of GPU memory for training and 6GB for inference. Utilizing a caching mechanism, we efficiently completed 35 training sessions and over a hundred inferences, each training session taking around 4 hours for 1000 epochs.
4.3 Performance of MoG on MQA Task
As mentioned above, we test the effectiveness of MoG on MQA datasets. For each question, the RAG system is tasked with choosing the best answer(s) from the given options. The performance of the entire RAG system is measured by the Exact Matching accuracy of the answers. To prevent knowledge leakage, only the question is used (options excluded) to retrieve reference documents from the external knowledge database. The router of MoG guides the retrieval system to choose the optimal granularity. Qualitatively speaking, when tested on different datasets, the router trained with Textbook corpus shows a preference for different granularity levels. For instance, on the PubMedQA dataset, the finest granularity snippets are selected most frequently. This is because the questions in PubMedQA are typically precise and can be answered with short reference snippets. An example result figure is shown in Figure 4, with a more detailed discussion included in Appendix D.
@@ 4.3 Performance of MoG on MQA Task
Table 1: Accuracy of Medical Question-Answering task with MoG (trained with MedCorp), best results in bold.
MoG is integrated into one same RAG system, with which we conduct the MQA task. Backbone LLMs are altered to cover some of the popular ones, such as ChatGPT Brown et al. (2020), InternLM2 Cai et al. (2024), Llama3 AI (2024), GLM3 Du et al. (2022), and Qwen1.5 Bai et al. (2023). The router is trained with the soft labels built using RoBERTa Liu et al. (2019), this choice is justified by the experiment in Appendix E. To investigate the effect of varying the number of candidate snippets on RAG system performance, we conducted experiments with different snippet counts (details are provided in Appendix F). The retriever is fixed as BM25 Robertson and Zaragoza (2009), with a further discussion on the performance of different retrievers included in Appendix G. In Table 1 we present the results obtained with the router trained with MedCorp corpus (the results obtained with routers trained on four single corpora are presented in Appendix H).The results are compared with two baselines: CoT and MedRAG. CoT baseline adopts Chain-of-Thought Wei et al. (2022) prompting and does not leverage any external knowledge database to enhance its response. MedRAG baseline is a simple RAG system with only 1 candidate granularity level introduced in MedRAG paper Xiong et al. (2024).
The results demonstrate that MoG consistently enhanced the performance of the RAG system across different backbone models when compared with MedRAG, though not necessarily better than CoT. The reason is that the RAG system we used has no noise filters or any quality control mechanism, thus the noise is injected along with the knowledge via prompts. A detailed analysis of the number of samples improved or degraded by the application of MoG is included in the Appendix I, in which we manually verified that the majority of degradation is caused by noise. We also find that MoG improves the accuracy score more when applied on weaker LLMs (like ChatGLM and Qwen), probably because they have less knowledge stored in their internal parameters and, thus could benefit more from the retrieved snippets.
@@
Table 2: Accuracy of Medical Question-Answering task with MoGG (trained with Textbooks), best results in bold.
4.4 Performance of MoGG on MQA Task
Similarly to the experiment of MoG, we test the performance of MoGG on MQA datasets. We tested only the performance of MoGG with routers trained on Textbooks and StatPearls corpora because training with the other two much larger corpora is too time-consuming. In Table 2 we present the results obtained when trained with Textbook corpus (results obtained with StatPearls showed in Appendix J with similar patterns). From the table, we can tell that MoGG can further improve the averaged accuracy scores. By comparing Table 1 and Table 2, we can find that, even when trained with significantly fewer samples (Textbooks corpus is a tiny subset of MedCorp corpus, accounting for only about 0.2% of all the snippets in MedCorp), MoGG brings more significant improvement in terms of the averaged accuracy score with respect to MedRAG than MoG. This finding highlights that MoGG is more efficient than MoG thanks to its flexible way of organizing the reference snippets (in the form of a graph). We conducted an ablation test with Llama 3 and the results are included in Appendix K.
There is a general performance drop in the metrics in Table 2 compared to Tabel 1 because the Textbooks corpus is only a tiny subset of the MedCorp used in Table 1). To facilitate the comparison, we conducted the same experiment in Table 1 with only the Textbooks corpus, and the detailed results are reported in Appendix L.
4.5 Execution time and storage efficiency
MoG(G) is proposed to increase the precision and recall of the retrieval phase at a reasonable cost of computational efficiency because the quality of the retrieved chunks is prioritized for application scenarios like the medical environment. We measured the average inference time with the different number of candidate granularity levels, and the results show that increasing the number of granularity levels will only increase a marginal increase in execution time. (Details in Appendix M)
In terms of storage, while additional space is required to store the embeddings of the corpus at different granularities, we only need to store one copy of the corpus and five sets of embeddings. This engineering optimization results in a space requirement of only 2.7 times the size of the original corpus. We believe this represents an acceptable overhead. Furthermore, in our RAG system specialized in the medical domain, a corpus containing about 10GB of plain text can already cover a wide range of questions. Thus, we believe this will not be an obstacle to MoG(G)’s wider application.
We believe this overhead of computing resources added is worthwhile because in most cases, MoG performs much better than MedRAG and CoT baselines. On smaller models, MoG shows an average improvement of 5% compared to MedRAG and 8.7% compared to CoT. Given the dataset size of approximately 7000, these improvements are statistically and practically significant.
@@ 5 Limitations and Broader Impacts
Our work serves as an early trial in dynamic chunking strategy, with the following major directions for improvement. (1) MoG(G)’s candidate granularity levels are manually assigned. It could be more efficient if an algorithm automatically set these granularity levels to avoid excessive grid-searching for parameter optimization. (2) Current router uses only the semantic information of the input query to predict the best granularity level. (3) Previous studies have demonstrated that the use of length normalization is crucial in information retrieval-related fields. This paper primarily focuses on applying the concept of Mix-of-Experts to this application scenario. In future work, we will also take related issues into consideration. Incorporating more information (like query type or expected response length) into the router can potentially improve the results. However, the router also introduces a new security risk: a compromised router could redirect knowledge retrieval to malicious sources, injecting incorrect or even harmful information into the backbone LLM. Therefore, it is crucial to protect and monitor the router to mitigate this risk.
@@ 6 Conclusion
In this work, we present MoG, a mechanism to dynamically choose the best granularity when retrieving information from an external knowledge database. When applied to a RAG system, MoG helps retrieve more relevant information while reducing noise. MoG is further extended as MoGG, where reference documents are pre-processed as graphs. This extension allows distantly situated information to be retrieved simultaneously, overcoming the limitations of a fixed top-$k$ selection strategy. Finally, we introduce a soft label guided loss function to address the difficulty of backward propagation with top-$k$ selection, which could benefit future research.
Acknowledgements
We would like to express our gratitude to all the supervisors and colleagues at Shanghai Artificial Intelligence Laboratory for their invaluable insights, feedback, and support throughout the research process. We also thank the reviewers for their constructive comments, which greatly improved the quality of this paper.
Appendix A More Discussion About Selection Process
The intuitive selection process would be to select the top snippets with the highest weighted similarity scores. However, this method is not adopted for the following reasons:
(1) It is biased, because more-coarsed snippets tend to have higher similarity scores, even though they might include more noise.
(2) It is not rebust to rely on the router’s output weight to control such imbalance between granularity levels, because the router is not traine for this objective. The router is trained with soft labels, which are approximate training signals that can only teach the model to distinguish the 1st and 2nd optimal granularity level from the rest.
(3) In practice, when $\mathcal{K}$ is too large, it has to be stored in separated vector databases. A common solution is use one vectorbase to store one granularity. This solution will make the similarity scores across granularity levels NOT directly comparable with each other.
Based on these reasons, we designed the selection strategy presented in the paper. The rationale behind the proposed strategy is that: Since we aim to extract the globally optimal snippet, we need a way to calculate a “global highest similarity score”. Given that the finest-grained snippets are the only ones that appear across all granularity levels following our initial chunking approach (fixed-size, non-overlapping), we calculate the weighted sum based on these finest-grained snippets.
Appendix B Details of QA Datasets
B.1 MMLU-Med
The Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al. (2021) evaluates the multitask learning capability of language models. While the full MMLU dataset encompasses 57 different tasks, we specifically extracted the medical questions for our tests, totaling 1089 questions.
B.2 MedQA-US
MedQA Zhang et al. (2018) is a multiple-choice QA dataset derived from professional medical board exams. It is available in Simplified Chinese, Traditional Chinese, and English. For our experiments, we used 1273 questions from the English version.
B.3 MedMCQA
MedMCQA Pal et al. (2022) comprises a large number of questions from the Indian medical entrance exam, covering 2400 healthcare topics and 21 medical subjects. For the MIRAGE benchmark, we utilized the “dev” set of the original MedMCQA dataset.
B.4 PubMedQA*